>
Home/Richard Williams

About Richard Williams

This author has not yet filled in any details.
So far Richard Williams has created 27 blog entries.

Apiary Record: 25th September 2016

Putting the colonies to bed for the winter, checking that they have enough stores and that entrances are down to a minimum.   Still a bit too early for mouse guards and woodpecker protection.    Removal of Varroa treatment and varroa boards also undertaken.
Colony #6   was on a National deep, but we combined Colony #1, which was on a 14×12 several weeks ago.   We had put the 14×12 under the deep after the two colonies were united, to encourage  them to take up the stores in it,  but they thought  otherwise and pulled out the entrance block, gaining access to the main entrance and so filled it with stores rather than taking it up into the National Deep. We went through the National Deep and used any stores left to help one of the Nuc’s and then removed it, leaving just the 14×12 and a full super on top.
We also decided to under-super the colonies,   putting full supers under the brood box’s, to help them get at the stores in the cold months to come.   Removing all the Queen excluders  to clean them ready for next season.
Weather was Sunny to start with  17 Degrees C ,  but became overcast with a shower of rain towards the end.  No wind to speak of.
Hive #2   Inspected by Richard  W.     Apiguard and Eke removed, top super 3/4 full and lots of stores in 14×12 Deep. Qx also removed.
Hive #3  Inspected by Richard W.     Apiguard and Eke removed, stores a bit light but incoming Balsam nectar and ivy could fill.  check next time..  may need feeding if weather bad.
Hive #4  Inspected by Richard W.  Apiguard and Eke removed.  Full super put under brood box and another full super put above a clearer board to give (bee free) to a nuc ( 16 B).
Hive #5  Inspected by Peter Reader,  Apiguard and Eke removed,  Full super put under brood box and Qx removed.
Hive #6  Inspected by Graham Rowden,  Apiguard and Eke removed, bees did not lift stores up into National Deep, so left in 14×12, ND removed, frames of stores given to 16 C.
Hive #7  Inspected by Richard W.   Apiguard and Eke removed. Qx removed and full super left on top as this has two ND brood boxes.
Hive #7A Inspected by Peter Reader. Apiguard removed with Eke. All Shallow boxes.  Qx removed,  top super full.
Hive #8  Inspected by Graham Rowden,   Apiguard and Eke removed.  Under-supered.
Hive #15F Inspected by Graham R,  Apiguard and Eke removed.
Nuc #16B  Inspected by Graham R.  Brood chamber – National Deep. Queen still laying nice brood pattern, BAS,  Full super from hive #4 added under deep.  colony to be used again for observation hive at Blackmoor.
Nuc #16C  Inspected by Graham R.   Brood chamber -National Deep, Light on stores, so given Deep frames from Hive#6,   Rapid feeder put on with 1 ltr sugar 1:1

Apiary Records: Tuesday 19th July

Was meant to be SBI  recheck on colonies to give clearance from EFB.   Due to SBI being off sick it was cancelled at the last minute,  but as we had delayed going through the colonies on Sunday we had to check them over.
Weather :-  HOT  and Sunny,  Little wind,  24 degrees.
Hive#1    not gone through as poss Virgin Queen
Hive#2    not gone through as poss virgin Queen
Hive#4    Inspected by Richard and Liz.   Brood Chamber:   National Deep.  QNS, no Qc’s. Brood and new eggs, room/stores ok.   4 supers,  3 ready to take off.  –  lots of play cups but none charged.  Temper 7
Hive#5   Inspected By Richard and Liz.    Brood chamber: Deep and Half.   QNS,   3 hatched Qc’s.  1 deep frame of brood/eggs . lot of drones – first time gone through for a month.  Temper 7
Hive#6  Not gone through as poss virgin Queen.
Hive#7  Not gone through
Hive# 7A    Inspected by  Richard and Liz,   Brood chamber:  4 Shallows ( Peter Street’s Method)                QNS,  Several play cups, not charged,  5 f bas in top super, room and stores ok. temper 8
Hive#8     Inspected by Richard and Liz.    Brood chamber: Deep.   QNS,  no Qc’s.  2 f BAS,  stores ok but room very limited,  Q laying right up to the frames .   2 supers, both near full, needs another next visit.  Temper  9.
Hive# 15 F,   Inspected by Richard and Liz,  Brood chamber: Deep.    QS,   1 QC,  Knocked down. 6 f BAS,  stores /room ok.  1 super( nearly full) –  Needs another Super next visit. Temper 9
Hive#15 H,    Inspected by Richard and Liz,  Brood chamber: Deep.   QNS,  3 Charged Qc’s knocked down. 6 F BAS with new eggs.  lack of room to lay so made up another Deep and moved another 3 ft towards Hive 3 slot.   Temper: 9.
Nuc’s  16B and 16C  not gone through.

Apiary Record : 10th July 2016.

Still under EFB shutdown until being rechecked by the SBI on the 19th July.  So quick check through, making sure no colonies were about to swarm and that they had enough room,stores and that they had no sign of any disease.
Weather:   Sunny after earlier rain/drizzle.  19 degrees C,  Breezy.
Hive#1   Inspected by Liz,    Brood Chamber  14 x 12,  QNS,  No Qc’s,  No Brood, No eggs, stores/room ok.   1 super  75% full,   Queenless..  Introduced a frame with 2 QC’s from 15H( soon to be the new Colony 3).  Drone Trap removed,  far too many drones in the hive. leave 4 weeks.  Temper : 9
Hive#2   Not Checked as sealed QC left on f2 on 3/7/16    will be checked on 19/7 by SBI
Hive#4  Inspected by Liz.   Brood chamber -National Deep.  QNS,  2 Qc’s Brood  AS,  and new day old eggs,  2 Supers  nearly full, stores/room ok.   –  Knocked down both charged Qc’s.  Temper: 7
Hive#5  Not gone through as poss virgin Queen.
Hive#6  Inspected by Liz,    Brood chamber,  Deep and a Shallow.   QNS, possibly swarmed.        4 charged Qc ‘s   Brood  5D, 4S,  2 supers, top full, –  knocked down 2 Qc’s leaving 1 sealed, and 1 charged but open, on nframe with red pin.  No eggs seen at all.-  leave for 4 weeks.                   Temper:  6.5
Hive#7  Inspected by Liz,   Brood chamber: Deep.   QNS,  no Qc’s, stores/room ok,  2 supers.     new eggs in drone comb only, so may be a drone laying queen, Combined with Nuc 16A as 2 Sealed Qc’s on f5.
Hive#7A   Inspected by Richard,    Brood chamber:  4 Shallows.   QNS, no Qc’s.  7 frames of brood, all in top super. some 6 -7 day old larva,  no new eggs.  Q gone?  Dirty comb in bottom supers needs replacing.
Hive#8    Inspected by Richard,    Brood chamber : Deep.   QNS, 2 charged Qc’s,  6 f BAS, new eggs, 2 supers, top full.  stores/room ok,   Charged Qc’s taken down.  lots of bees looks like they may swarm.
Hive#15F  Inspected by Liz.     Brood chamber: Deep.  QS,  no Qc’s.  6 f BAS,  stores/room ok,     1 super,  –  Knocked down several play cups.
Hive#15H   Inspected by Richard.   Brood chamber: Deep.   QS,  2 QC’s ,  7 f BAS and new eggs.  Stores/room ok,  1 super . Part filled,-   Frame with 2 Qc’s given to Hive 1.    Temper:  9
Nuc # 16A   Combined with Hive#7.   but no Q found, 2 sealed QC’s  left.
Nuc# 16B   Not gone through – though bees bringing in pollen
Nuc# 16C  Not gone through –  bees flying
 
 

Apiary Record: 26th June. From memory….

To follow the EFB  “lockdown” to the letter, we had to leave all the record sheets in the container so a rough guide from memory (known to be unreliable ).
Hive #1   Not checked as poss Qc / Virgin Queen.
Hive #2  Checked and QS,   BAS with new eggs.   lots of play cups taken down but none charged. top super full ready to take off when we can.
Hive #3   Culled
Hive#4   Checked and all  OK.   Top 2 supers full,  3rd filling.
Hive#5   Not checked  as  poss Virgin Queen.
Hive #6  Checked and QS ,  BAS and eggs,   6 Qc’s  some sealed,  another Nuc made up from a frame with two nice Qc’s on , along with a frame of brood and a frame of stores.  The remaining Qc’s were taken down.  Very nice tempered colony.
Hive #7  Not Checked as poss VQ.
Hive #7A  Not checked as poss VQ.
Hive #8  Not checked as poss VQ.
Hive #9  Warre Hive   – Culled as a precaution, due to being a very small colony and thought to be more susceptible to EFB and not making up.
Hive#10  Top Bar Hive –  Also a very small colony and culled as per Warre hive.
Hive #15F  Ex  Nuc, now in Deep, with super added.  building up nicely.
Hive #15H  Ex Nuc,  also building up, QS. a few playcups, 5f BAS , 1 super – moved another meter towards No3 space.
Nuc #16A   From Hive 4,  QNS,  2 play cups, waxmoth larva killed. making up nicely, 4f BAS lots of eggs.  Temper 10.   2 varroa in 4 weeks.
Nuc #16B  From Hive 4, 19/6  left as poss VQ
Nuc #16C  From Hive 6  26/6  put into wooden Nuc, furthest from Apiary on stand in ferns,

Apiary Report: 19th June 2016.

The Apiary still on EFB lockdown.   But the colonies still need a go through, and the grass needs cutting again.    Two custodians ,assisted by David Parkinson went through the colonies to check for Queen cells and any signs of further problems with EFB.  Temperature 18 degrees, Cloudy.
Hive# 1  Not gone through as 2 Qc’s left on f7 on 12/6,  no need to check till 3/7.
Hive#2  Inspected by David P.   Brood Chamber 14 x 12.   QS,  8 Qc’s taken down as unable to make up Nuc with 14×12 frames.  9 f brood BAS and eggs, stores ok.   Temper 8.   Super added above brood box as very little space for Q to lay up.
Hive#4  Inspected by Liz,    Brood chamber : National Deep.  QS, 12 play cups taken down,(not charged).  BAS and eggs,  room and stores ok.  3 Supers,  Top Full. next , threequarters full, bottom filling but less in than last week as being used during bad weather. Temper 7.
Hive#5   Inspected by  Richard W,    Brood chamber: Deep and a half,  QNS,  No eggs but a few sealed brood left, No Qc’s.   2 Qc’s from Hive 6, on frame with brood – no bees, put in. (nice calm bees).   Leave Till  10/7..    Temper 4.
Hive#6   Inspected by David P.    Brood chamber : Deep and a Half,   QS,  6 Qc’s,  BAS and new eggs.  stores and room OK.  Temper 9.    – moved 1 frame with 2 Qc’s on into Hive 5. Other Qc’s taken down.  Replaced with frame of foundation, giving more room in brood nest.
Hive#7  Inspected by Richard W.   Brood Chamber: Deep.   QNS,  3 open hatched Qc’s.  no brood or eggs. ( poss virgin Q ), stores room OK, 2 Supers, Top full.   Temper 9.  bees very calm so poss Virgin Q present.  Put back together ASAP.
Hive#7A  Inspected by Richard W.   Brood Chamber 3 Shallows.    QNS.  3 open Hatched Qc’s. but 4 shallow frames of 4 – 5 day old brood, and some eggs. so new Q now laying up.  Stores and room OK,  Bottom super MT, so put on top of QX.     Temper 9.
Hive#8    Not gone through as Qc’s left on 12/6.   – Check again  3/7.
Hive#10   Top Bar Hive.     Inspected by Liz/Richard,     QS,  No Qc’s,   4 f BAS. and eggs. 2 stores room OK.   Temper 10.  – Ants still a problem nesting between the top bars.   Must bring a hoover to clear hive debris from OMF.!!!
Hive# 15 F   Inspected by Liz/Richard.   Brood Chamber: Deep,  QNS, no Qc’s,  5 f  BAS and 1 f eggs,  3 stores 2 undrawn  f for space.  added super above QX for added space. Temper 10.
Hive#15 H  Inspected By Liz.   Brood Chamber: Deep.  QNS, few play cups.  5 f BAS, stores and room OK,  1 Super starting to fill.  Temper 7,    Moved 3 Feet towards space at no3, After Inspection, but bees still a bit “grotty”.( Liz’s term)
Nucs 16A and 16B  not gone through,  but area of ferns cleared from around Nuc’s.  Bees Flying.
 

Apiary Records: 12th June.

Under Lockdown restrictions after EFB confirmed two weeks ago.   Weather  20 degrees, cloudy.
Hive#1  Inspected by David Parkinson.   Brood Chamber: 14 x 12.
QNS,   Lots of Qc’s after killing Q last week, all taken down bar two best ones on frame 7.              6 f brood,   3 f stores,  room ok,  1 super most empty.  Temper 8.
Hive#2  Inspected by Liz.    Brood Chamber: 14 x 12.
QS,  8 play cups ,not charged,  8 f brood, stores and room OK. 1 super filling. Temper 7,   Not yet pulling shallow varroa trap.  3 stings, not happy bees.
Hive#4  Inspected by Gaye.    Brood chamber:  National Deep.
QNS,   1 torn down Qc,  8 f brood BAS,  stores and room ok,  3 supers, 2 full.  Temper 8.   Swapped  some of the full super frames for empty with Hive#1.
Hive#5  Inspected by David,   Brood Chamber:  Deep and  a Half.
QNS,  6 Qc’s,   4 f sealed brood, no eggs, some larva, stores and room ok.  2 supers. Removed all but 2 Qc’s as poss queenless.  1 frame with 2 Qc’s given to Hive#8.   Very aggressive bees, Temper 4.  Lots of stings.( poss due to being Queenless)
Hive#6   Inspected by Richard,  Brood Chamber: Deep and a half.
QS,  No Qc’s, 8d f brood,  6s f brood all stages, stores and room ok,  1 super nearly full. Temper 9
Hive#7   Left alone as poss Virgin queen,   ( Qc’s left 22/5)  check next week.
Hive#7A  Not inspected.
Hive#8  Inspected by Richard.   Brood Chamber: Deep.
No Queen,  No Qc’s.  Little brood mainly Drone left. Lots of stores and room.  Added frame of brood with 2 Qc’s, 1 open 1 sealed, from Hive#5 ( no bees).   Bees very agitated. Temper 6.
Hive#10  Inspected by Liz,    Top Bar Hive.
QNS,  no Qc’s,  3 f BAS.   2 f stores, room ok.   Temper 10,   Wax moth still evident and ants a problem.
Hive#15F   Inspected by Gaye,  Nuc in National Deep.
QNS, no Qc’s,  6 f Brood . BAS.  1.5 f stores, room ok, No super.   Temper 10.
Hive#15H  Inspected by Liz.    Nuc in Deep.
QNS, no Qc’s,   5 f BAS and 1 day old eggs,  stores and room ok,  1 super.   Moved bees 3 ft towards  no 3 position,  before inspection, so bees very agitated, better to move after inspection.
Nuc’s   16 A and 16 B   (both made up from colony 4) Not gone through but bees flying ok.
 
 

Apiary Records : 5th June 2016

First visit after lockdown.   The custodians and Rowan, attended to go through the sterilization of all apiary equipment that was likely to come into contact with bees from colony 3 and to burn and clean equipment from the Warre hive that was culled due to the possibility that a weak colony could easily catch EFB.  Rowan did a grand job with the fire ( day off for a firefighter!).
Temperature 25 degrees C  and Sunny.   No wind.    –  A lovely day.
Hive#1:   Inspected by Richard.  Brood Chamber:  14 x 12.
Queen found at last and Killed.(Regicide), due to high levels of chalk brood.  No Qc’s.   6 f BAS, stores and room ok, 1 super starting to fill,     Check next week and reduce Qc’s to 2. Temper 9.
Hive#2:    Inspected by Richard.   Brood chamber   14 x 12.
Queen seen,  2 qc’s taken down,   8 f BAS,  stores and room OK,  1 super Part filled. temper 9.  Bald brood still evident. signs of varroa damage.  big colony.
Hive#3:    Culled due to EFB.
Hive#4:   Inspected by Richard,  Brood chamber: National Deep.
QNS,  5 sealed Qc’s.  7 f BAS and day old eggs. Stores and room ok,   3 supers, 1 full, 1 new foundation, bottom raised up from brood chamber with brood on, as advised by the SBI, as mainly stores.    Frame with Qc’s on used to make up another Nuc, along with a shallow frame of stores. To be replaced ASAP.  Deep frame replaced with foundation to give more room in brood nest.   Other Qc’s taken down.   Temper 9.
Hive#5:   Inspected by Liz.   Brood Chamber: Deep and a Half.
QNS,   10 Qc’s  play cups, some charged,  taken down .  6f BAS, stores and room OK. 2 supers.  lots of happy bees.   Temper 8.
Hive#6:  Inspected   by Liz.   Brood chamber: Brood and a half.
QNS,   8 Qc’s,  some play cups, taken down. 8 f BAS, +day old eggs. stores mainly in shallow brood box.   1 super, part filled.  Temper 7.
Hive#7:  Inspected by Mel.     Brood Chamber:   Deep.
QNS, no Qc’s.  no brood,  stores and room OK,  2 Supers , 1 Full.   No evidence of eggs yet.               This was an artificial swarm ( 18/5 ) .  with Q left on foundation and fed syrup.  is she still there?
Hive#7A:  Inspected by Mel.    Brood chamber:  4 Shallows.
QNS, no Qc’s.  some brood still but no new eggs. stores and room OK.   no supers.  Temper 7. bees very agitated.  Should have been a queen cell somewhere !
Hive#8:   Inspected by Mel.   Brood Chamber : Deep.
QNS,  12 QC’s lots of brood but no eggs.  ( Q killed 29/5) stores and room OK.  1 super.
Hive#9:   Culled on the advice of SBI
Hive#10   Top Bar,   Not Inspected
Hive#15F  Inspected by Liz,    Brood Chamber: Deep.
QNS,  No Qc’s,   4.5 f  BAS,  stores and room ok.  No Supers.  temper 9.   some play cups. bees on frames 6 and 7,   It was hoped to use this colony in the observation hive for the coming two fete’s but due to the lock down we are unable.   We will now make up to a full colony.
Hive#15H     Not Inspected.
Hive#16A     NUC made up last week with Queen cell, not inspected, but bees still flying a good sign.
 
 
 

Apiary Report: 29th May. Health Check by SBI – the dreaded EFB confirmed

All was going so well , starting from the furthest colonies from the entrance, Top bar, Warre, No# 8, ect until we got to No3 .  All went pear shaped.     Debbie Burton ,the Seasonal Bee Inspector, discovered a few cells of EFB.  This colony had been small from the start of the season and slow to build up,  Debbie thought that the EFB may have been present over winter and not shown itself until now.    For those present it was a very useful experience to see the signs of EFB at first hand and what to look for in the brood frames.    If proof were needed, do not try to keep a failing colony going and certainly do not combine it with another colony unless you are sure that it has no disease.   Once condemned ,the colony had to be culled ASAP later that evening, when as many flying bees had returned as possible and the frames and bees burnt later.    Debbie advised that the boxes, floor, Qx, crown board  and roof would need to be scorched with a burner and removed to the container, for possible use next year.    A very sad end to the day, but necessary for the health of the other colonies.
Temperature  20 degrees and sunny.  Little wind.  All hives inspected by Debbie Burton  SBI  working from 15H back to No1
Hive#1   Brood chamber:     14 x 12,   QNS,  no Qc’s,   5 f  BAS,   stores ok, room ok. 1 super.   temper 10.   It was hoped to find the Queen and kill, as the colony is suffering very badly with chalk brood. – Next time…
Hive#2    Brood Chamber:    14 x 12,    QNS,  no Qc’s,  8f BAS,     stores ok, room ok,  1 super,  temper 8.  Wax moth grub found and killed, some bald brood evident, but Q laying well.
Hive#3    Brood Chamber: National Deep.  QNS, 3 f  BAS, – Ants present in hive, bees “jittery”, poss chronic bee paralysis,  then grey misshapen larva discovered, tested for EFB and confirmed positive.  B****r.   To Cull bees after flying bees return by sealing hive up and dosing with petrol.   Frames and bees to be burnt later. Apiary now on lockdown,  Nothing to be removed from site.
Hive#4    Brood Chamber:  Deep and Half,   QS,  9 Qc’s,  12 f BAS,  stores ok, room ok, 2 supers.   Some Qc’s sealed but Q still present.  Frame of brood with 2 Qc’s and f- food put into poly Nuc,with some extra bees to make up a new colony.  Put on stand in fern area away from trees.  Remaining  Qc’s taken down, and removed frames replaced with foundation to give more room.  Super frame with lots of drone cells removed for varroa check, replaced with DC.
Hive#5   Brood Chamber:  Deep and Half.    QS,   2 Qc’s one sealed.  needed more room so moved end frames into brood nest,  good distribution of new brood in shallow , put stores  up into Super to give more room in Shallow.
Hive#6    Brood Chamber:   Deep and Half,   QNS.   no Qc’s,   8 f BAS,  stores ok, needs more room, day old eggs, but possible ready to swarm, moved shallow f, into brood nest to give more room.  Wax  moth present in f 5 and 8.
Hive#7    Brood Chamber:   New Deep.  QNS,  12 charged Qc’s,   No brood, stores ok, room ok.  Artificial swarm made on 18/5.  but Q may have gone.  2 Qc’s left, others taken down.  New super of DC added as Super full.
Hive#7 A  Brood Chamber:   4 Shallows,  QNS,  12 Qc’s     8 f BAS,  stores ok, room ok. 1 super added.   Ants coming in.  1 sealed Qc -small, opened and larva dead. all brood sealed. QC opened so may be virgin Q ,  other Qc’s taken down.
Hive#8    Brood Chamber:   National Deep.  QS,  7 Qc’s,   8 f BAS,   stores ok, needs room,  Qc’s charged,  drones being laid in worker cells, Chalk brood present, sbi advised Regicide (Q killed)  widespread pepperpot brood. sealed QC left on F6.
Hive#9    Warre Hive.    2 brood chambers,  QS,  no Qc’s, 2 f BAS,  not much stores, small colony, old comb needs replacing if we are to keep going with colony.  No sign of any disease,   few ants around.  Solid floor may have caused mouldy frames, poss due to lack of ventilation over winter.
Hive#10    Top Bar Hive.   QNS, no Qc’s,  3 f BAS, little stores, room ok.     Wax moth and bald brood present.   one cell found with 2 eggs in!  Building but still weak. – ( needs swarm dumped in , as per Peter Reader)
Hive#15F   Brood Chamber:  National Deep. QS,  no Qc’s,  4 f BAS,   Sac brood (early stages) present on f6. some bald brood also present, along with varroa damage.
Hive#15H   Brood Chamber:  National Deep,   QS,  no Qc’s,  4 f BAS,  varroa damage, Sac brood and Bald brood present.  Laying drones in worker cells. stores ok but short of room, building wild comb in roof above cb.  A super of DC was added to help.
 
 
 

Apiary Report: Sunday 15th May.

Another lovely Sunday with lots to do.  Now is the time that the colonies are getting going, and its the ideal time to get hands on experience on colony manipulations.    Lots of activity in the hives, and all bar one or two making up nicely.    Nectar and pollen incoming, thanks to the warm weather and the Queens laying up to nearly her maximum, ready for the main flows in early June.        Number one priority was to go through the Warre hive for the first time this year.   What a mess,  lots of unattatched comb,  and floor debris , but the bees were still there and surviving.  The problems of operating this type of hive were apparent and though mimicing the natural hollow tree, it is not easy to check through regularly and as for swarm controls, very difficult.
Temperature   16 degrees C,    Sunny periods.  little wind.
Hive#1    Inspected by Brian Herbert
Brood  Chamber:  14 x 12,    QS,  Play cups,  6f Brood at All Stages,  (BAS)  3f stores,  room ok.  1 super . empty. Temper 7,   vb out.  –   Lots of chalkbrood,  A shallow frame was put into broodnest for drone trapping part of IPM.       Recomended  to requeen ASAP  to overcome chalkbrood.
Hive#2    Inspected by Brain
Brood Chamber;  14 x 12,    QS,  No QC,   4.5  f BAS,  4 f stores, 2+ room,  1 super, Empty.  vc 2, taken out.  Temper 10 Frame with brood on found outside brood nest ,next to stores so moved in. Replaced part drawn frame with shallow for drone trapping.
Hive#3      Inspected by David Parkinson
Brood Chamber:    National Deep.   QNS, No QC,  only 2 f BAS, 1 f stores, room ok,  no supers. temper 10.                       Weak colony.  not building up much, Wax Moth found , may be best to Kill Q and Combine.
Hive#4   Inspected by  David
Brood Chamber:  National Deep and half (shallow).  QNS,  no QC,  8d + 3s f BAS, 3f stores, room ok, 2 supers temp 9   colony building nicely, Healthy hive full of bees.  vb 0. taken out.
Hive#5   Inspected by Anne Chantell Ballard.
Brood Chamber:    National Deep. Shallow added.    QS,  1 QC taken down, not sealed. 7 f BAS,  11 f stores, room ok.     1 super, Temper 8 .   –  cramped in deep so super added to brood chamber.  one deep frame replaced with shallow for drone trapping.   entrance block removed,  Consider an artificial swarm next week if more QC’s
Hive#6   Inspected by Anne
Brood Chamber:   Deep,   Shallow added.  QS,  no QC.  6 f BAS, 3 fd stores,  1 super empty, Temper 9. live wax moth larva on vb,  vb removed  and entrance block removed.
Hive#7   Inspected By Liz Hawes
Brood Chamber:   4 Shallows,  QNS,   10 Qc’s. 2 left on.   20+ f BAS. stores ok. Q excluder  added  under new super of drawn comb,  some full frames of stores moved up from brood chamber into super. replaced by  f drawn comb. will have to do an artificial swarm ASAP.
Hive#8     Inspected By Liz and Mel
Brood Chamber: National Deep.  QS, No QC,  9 f BAS, 2 f stores, room ok, 2 supers. temper 9.  Checked drone brood for varroa, none found, entrance block and vb removed.
Hive#9    Inspected by Richard Williams
Warre Hive.       QS, No Qc’s. 2 f BAS, 1 f stores, room ok,  temper 10.   –  Taped back up several loose frames, very dark and bit mouldy but being used..   better new comb in bottom box.  solid floor cleaned of debris and entrance cleared of old foam.   Small colony.
Hive#10   Inspected by Peter Reader
Top Bar Hive.    QNS, no Qc,   2 sides of BAS,  2 stores,  room ok. temper 10. –  queen has sprung back into action with now 2 sides of brood, but still small colony.   making up slowly.
Nuc  15 F    Not gone through as brood in observation hive at Milland Show.
Nuc 15 H    Inspected by Richard
Brood Chamber: Deep.   QNS, No Qc’s. 3.5 f BAS.  1 f stores, room ok.  Temper 10. rapid feeder taken off,   Moved a further 3 ft, onto concrete platform.
 

Apiary Report: Sunday 8th May 2016. What a lovely day..

Hot and Sunny, what could go wrong!!!    How about lack of new members, where were you all?   Those that did attend very much enjoyed themselves, with lots of hands on experience opening hives and checking bees.    Things are hotting up in the colonies with most hives showing lots of stores coming in and increasing build up of brood.
Apiary Sundays are a very sociable afternoon, when we can exchange information on bee topics , problems with ones own hives, how to reduce varroa, what to look for in colonies affected with nosema, different hive set ups and general upkeep of hives.
Thirteen members attended Sunday, with only 6 of this years “New Bees”  .  Temperature  in the 20’s C  Sunny.
Hive#1   Inspected by Brian Herbert
Brood Box;    14 x 12     QNS, No QC’s , 4 f brood,BAS. 3 f stores, 2 f room,   1 super,  varroa count 5,  temper 8.              Empty super replaced with another as the frames were not sitting down below the box sides.  Several drone cells were lifted to check for varroa as a few bees were seen with DWV, ( Deformed Wing Virus) thought to be transmitted by high counts of varroa. None were seen on the samples taken.
Hive#2   Inspected by Brian,
Brood Box:    14 x 12 ,   QS,    No QC’s, 3.5 f Brood BAS,  2+ room,  Varroa count 30,  Temper 10.    Dusted with icing sugar for high varroa count.  Overcoming chalk brood problems last week but some diarrhoea noticed on super frames.
Hive#3   Inspected by  Peter Reader.
Brood Box:  National Deep.   QNS,   no QC.   2 f BAS  1.5 f stores,  no supers.  varroa count, Zero. Temper 10                      Still very small colony but building slowly.  good pattern of brood and stores ok. rapid feeder removed.
Hive#4   Inspected by Peter Reader.
Brood Box:    National Deep and Added Shallow.  QS,  no QC,  7 f BAS,  3f stores,   1 super.   no varroa, Temper 9.           Not much room left in deep, so added shallow of DC ( Drawn Comb), big strong colony.
Hive#5    Inspected  by Richard Williams
Brood Box:  National Deep,  QS,  no QC,  5.5 f BAS, 4 f Stores, room ok,  varroa count  3.  Temper 10.                                 Queen located and marked blue as last years. nice and fat still.  lots of stores coming in but not capped yet.
Hive#6   Inspected  by Richard Williams
Brood Box:  National Deep.    QS,  no QC,  5 f BAS, 3 f Stores, room ok,   1 super started filling, no varroa, temper 10.     Some chalk brood on vb,  no mouse droppings as previous weeks, small colony but building up ok.
Hive#7   Inspected by Liz Hawes
Brood Box:  3 All Shallows.( Peter Street)  QNS,  2 QC taken down as damaged, 20 f Brood,  35 f stores, still room ok.   high varroa count at 45.   Temper 9.     Open Queen cells damaged with removal of shallow box as on bottom edge. Lots of bees, strong colony, but bit cramped in shallows.
Hive#8   Inspected by Liz Hawes
Brood Box:  National Deep.   QS,  no QC,  7 f BAS,  stores ok, room ok,  2 supers.  varroa count 2.  Temper 9.                    brood that was in super has now all gone and being filled with stores.
Hive#9     Warre Hive –  Not gone through –  Priority for next week.
Hive#10   Inspected by Peter Reader  Top Bar Hive  Expert..
QS,  no QC, 1 f BAS, stores ok, room ok,    v count , 1.  temper 10.   Moved colony to be central, not making up very well.  same as last week.  50:50 sugar syrup in rapid feeder not taken down, now mouldy so taken off.
Hive#15F    Inspected by Richard
Brood chamber : National deep,  QNS,  no QC.  3.5 f BAS,  1 f stores,  room ok,  no varroa, temper 10.   Rapid feeder taken off ,  making up slowly.
Hive#15H   Inspected by Richard
Poly Nuc, now transfered into National Deep.   QNS, No QC,  3.5 f BAS,  2 f stores,  VC = 2 .  Making up nicely. will be used for display purposes soon,  Moved  3 feet towards concrete slab